OpenStudy (hiroromon):

How many moles are in the container? If a gas is compressed in a 5.00L container under a pressure of 25.0atm and at the temperature of 20.0 degrees C. R=0.0821L atm/mol K

OpenStudy (chmvijay):

PV=nRT n=PV/RT v=volume T=temperature in kelvin P=presure R=0.0821 n= no of moles can you find n by putting all these values in the above equation

OpenStudy (hiroromon):

i can try.

OpenStudy (chmvijay):

Ok good try it :)

OpenStudy (frostbite):

Remember to add to the answer that it is a approximation as we assume the gas is ideal :)

OpenStudy (hiroromon):

n = 25.0atm x 5.00L/0.0821 x 293 k ^ do i input that in the calculator?

OpenStudy (chmvijay):

yaaa LOL you can

OpenStudy (hiroromon):

erm i got a huge number . . . .

OpenStudy (chmvijay):

25*5 =125 right 0.0821 * 293 =24 right now u divide thme

OpenStudy (hiroromon):

oh

OpenStudy (chmvijay):

what oh

OpenStudy (hiroromon):

i got it this time

OpenStudy (chmvijay):

how much value you got LOL

OpenStudy (chmvijay):

yaaa may be @Frostbite

OpenStudy (hiroromon):

yeah i did . . . .

OpenStudy (hiroromon):

oops

OpenStudy (hiroromon):

thanks :)

OpenStudy (frostbite):

Correction: \[n \neq \frac{ pV }{ R }T\]

OpenStudy (chmvijay):

LOL tell me the answer :)

OpenStudy (hiroromon):

5.20 moles due to sig figs

OpenStudy (chmvijay):

yaaa ur right :)

OpenStudy (hiroromon):

weeeeeee

OpenStudy (chmvijay):

AWWWWWWWWWWWWWW